OpenStudy (anonymous):

f h(x) = 2x + 3, find the value of x if h(x) = 9. 21?

OpenStudy (anonymous):

yeah its 21

OpenStudy (anonymous):

No. Again, you got to flip the thinking. h(x) = 9 2x + 3 = 9 2x = 6 x = 6/2 x = 3

OpenStudy (anonymous):

Oh!! oh!

OpenStudy (precal):

I believe this problem was already posted and solve earlier today

OpenStudy (anonymous):

@annas Stop giving the same bad answers. I told you last time you were wrong and you come and do the same thing.

OpenStudy (anonymous):

@GT i dont know wht your talking about i m telling that she is right thats wht she ask for i m not solving questions and show me some respect boy..... you told me nothing

hero (hero):

@GT is correct: h(x) = h(x) 2x + 3 = 9 2x = 6 x = 3

OpenStudy (anonymous):

@annas how is she right with the answer of 21?

OpenStudy (anonymous):

ok its 12

hero (hero):

@annas , you might want to read the code of conduct

OpenStudy (anonymous):

i ve read it ...